# LLM-OptiRA: LLM-Driven Optimization of Resource Allocation for Non-Convex Problems in Wireless Communications

### v1

@@ Abstract
Index Terms:
Non-convex optimization, large language models, resource allocation, wireless communications
@@ I Introduction
Our implementation is openly available for research here: https://github.com/Bearrr310/LLM-OptiRA
@@ II-B Mathematical Model Construction
Next, the modeling of the constraints is conducted, for which the model needs to recognize the types of variables and the constraints. First, the model identifies the type of each variable $t_{i}$ (e,g. integer or continuous), forming a set of extracted optimization variables =$\mathcal{E}=\{(x_{i},t_{i})\}$. Next, the model identifies the numerical values $c_{j}$ related to the constraints and then analyzes the text to determine which optimization variable $x_{i}$ corresponds to each $c_{j}$. Following this, the model extracts the sentences $r_{ij}$ that mention these constraint relationships. Finally, these correspondences are placed into the set of constraint relationships $\mathcal{R}_{c}$, defined as
$\mathcal{R}_{c}=\bigcup\limits_{i,j\in\mathcal{E}}\{(x_{i},c_{j},r_{ij})\}.$ (1)
@@ II-F Feasibility Domain Correction
$\mathbf{y}^{(FDC)}_{l}=\begin{cases}\mathcal{F}_{adj}(\mathbf{x}_{0},\Delta\mathbf{x},\gamma)&\text{if }\mathcal{V}=0\text{ and }l<\left\lfloor\frac{L}{2}\right\rfloor,\\ \mathcal{F}_{re}(\mathcal{P},l)&\text{if }\mathcal{V}=0\text{ and }\left\lfloor\frac{L}{2}\right\rfloor\leq l\leq L,\end{cases}$ (7)
@@ III-A Dataset
The dataset is created by extracting 50 resource allocation problems from textbooks, research papers, and Matlab documentation[2][16][17]. These problems provide communication background descriptions, optimization formulas, and simulation parameters, forming the base dataset. Meanwhile, we utilize the LLM for data augmentation, enabling it to learn the core characteristics of these problems and generate 50 additional similar optimization problems in different contexts.
@@ III-C Evaluation Metrics
    Success $\text{Success Rate}=\frac{1}{|\mathcal{D}|}\sum_{\mathcal{P}\in\mathcal{D}}\left(\frac{1}{N}\sum_{i=1}^{N}\mathcal{V}_{\mathcal{P}}\right),$ (8)
@@
    Execution $\text{Execution Rate}=\frac{1}{|\mathcal{D}|}\sum_{\mathcal{P}\in\mathcal{D}}\left(\frac{1}{N}\sum_{i=1}^{N}\mathcal{Q}_{\mathcal{P}}\right),$ (9)
@@ III-F Maximum Iteration Count Analysis of ECL and FDC
ECL effectively increases success and execution rates in early iterations. In terms of success rate, ECL quickly identifies and corrects key errors in the early iterations, significantly enhancing overall performance. GPT-4’s success rate increases from 60% to 82%; GPT-3.5’s success rate rises from 25% to 58%, surpassing the PS + Refine baselines when $K=4$, indicating that ECL is effective for both high- and low-performance models. In terms of execution rate, ECL’s improvement mainly concentrates on the first few iterations, with GPT-4’s execution rate reaching to 96% by the fourth iteration. Although GPT-3.5 starts with a lower execution rate, ECL still significantly enhances the execution capability of this low-performance model, increasing it from 25% to 62%. Ultimately, the framework sets the maximum number of iterations $K=4$ to balance computational time and performance.
FDC demonstrates notable improvements in success rate and maintains steady execution rate enhancements with increasing iterations. For the success rate, GPT-4 rises from 19% to 81%, and GPT-3.5 improves from 6% to 52%, demonstrating FDC’s strong effectiveness in correcting infeasible solutions, while CoT and PS Refine do not show a similar increase. This improvement is due to FDC’s iterative - based adjustment mechanism. It corrects solutions during the initial optimization attempts and re - examines the problem when reaching $\left\lfloor\frac{L}{2}\right\rfloor$ iterations, effectively fixing infeasible solutions. In terms of execution rates, although FDC primarily focuses on ensuring solution feasibility, it also indirectly boosts execution rates. This consistent enhancement in feasibility supports the stability of the execution process, allowing FDC to optimize solutions even with a low initial execution rate. Ultimately, the framework sets five iterations $L=5$.

### v2

@@ I Introduction
Our implementation is openly available for research here: https://github.com/Tibbers0310/LLM-OptiRA
@@ II-B Mathematical Model Construction
Next, the modeling of the constraints is conducted, for which the model needs to recognize the types of variables and the constraints. First, the model identifies the type of each variable $t_{i}$ (e,g. integer or continuous), forming a set of extracted optimization variables $\mathcal{E}=\{(x_{i},t_{i})\}$. Next, the model identifies the numerical values $c_{j}$(where $c_{j}\in C$, $C$ is the set of constraint values) related to the constraints and then analyzes the text to determine which optimization variable $x_{i}$ corresponds to each $c_{j}$. Following this, the model extracts the sentences $r_{ij}$ that mention these constraint relationships. Finally, these correspondences are placed into the set of constraint relationships $\mathcal{R}_{c}$, defined as
$\mathcal{R}_{c}=\bigcup\limits_{x_{i}\in\mathcal{E},\,c_{j}\in C}\{(x_{i},c_{j},r_{ij})\}.$ (1)
@@ II-F Feasibility Domain Correction
[…]
@@ III-A Dataset
The dataset comprises 50 resource allocation problems extracted from textbooks, research papers, and Matlab documentation[2][16][17], covering communication backgrounds, optimization formulations, and simulation parameters. To expand the dataset, the LLM learns core patterns and generates 50 additional problems in diverse contexts.
@@ III-C Evaluation Metrics
    $\text{Success Rate}=\frac{1}{|\mathcal{D}|}\sum_{\mathcal{P}\in\mathcal{D}}\left(\frac{1}{N}\sum_{i=1}^{N}\mathcal{V}_{\mathcal{P}}\right),$ (8)
@@
    $\text{Execution Rate}=\frac{1}{|\mathcal{D}|}\sum_{\mathcal{P}\in\mathcal{D}}\left(\frac{1}{N}\sum_{i=1}^{N}\mathcal{Q}_{\mathcal{P}}\right),$ (9)
@@ III-F Maximum Iteration Count Analysis of ECL and FDC
ECL significantly improves success and execution rates, especially in early iterations. For success rate, ECL rapidly identifies and corrects key errors, boosting GPT-4’s success rate from 60% to 82% and GPT-3.5’s from 25% to 58%, surpassing PS+Refine when $K=4$, , indicating that ECL is effective for both high- and low-performance models. Execution gains are concentrated in initial iterations, with GPT-4 reaching 96% by the fourth iteration. Despite GPT-3.5’s lower baseline, ECL raises its execution rate from 25% to 62%. To balance performance and computation, the framework sets $K=4$.
FDC demonstrates notable improvements in success rate and maintains steady execution rate enhancements with increasing iterations. GPT-4’s success rate increases from 19% to 81%, and GPT-3.5’s from 6% to 52%, outperforming CoT and PS Refine. These gains stem from FDC’s iterative adjustment mechanism, which corrects solutions early and re-examines the problem at $\left\lfloor\frac{L}{2}\right\rfloor$ iterations to fix infeasibility. Although FDC primarily focuses on ensuring solution feasibility, FDC also indirectly boosts execution rates by stabilizing the solving process, even under low initial performance. The framework sets $L=5$ to balance correction depth and efficiency.
Acknowledgment
This work has been supported by the key research and development plan projects of Jiangsu Province under BE2022316 and Key R&D Program of the National Ministry of Science and Technology (2023YFB2905603)